A 5kg weight whose motion is constrained by vertically guides is dropped from rest 0.5m onto vertically aligned spring whose stiffness is 10kNm-¹. how fast is the weight travelling when the spring has been compressed by 50mm?


1
Expert's answer
2022-02-23T08:32:31-0500

"mgh=\frac{kx^2}2+\frac{mv^2}2,"

"v=\sqrt{2gh-\frac{kx^2}m}=2.2~\frac ms."